Project

General

Profile

Update 13 July

Status of running
Added by Kyberd, Paul over 2 years ago

The purpose of this update is to summarise the state of the experiment at the end of ISIS Cycle 2015/01a and to outline the priorities for the coming Cycle (2015/01b), which is about to start.

Status update:

By the end of 2015/01a, all of the MICE instrumentation had been commissioned and detailed monitoring had been made available in the control room. As an example, the first attachment shows a selection of the plots available from the Online reconstruction. Each of the detector systems is now represented.

Offline, substantial progress had also been made. MAUS now is able to reconstruct the data faster that we can record it; a really important step. The straight-track reconstruction too has been debugged and refined. As an example of the progress here, the second attachment shows the smoothed residuals from the track fit in each plane of both trackers. There are details to understand, but, the data and the reconstruction software can now be used to study the alignment, efficiency and calibration of the devices.

For the past week or so, the spectrometer solenoids have been cold, each containing a volume of liquid helium. The control system has maintained the magnets at a stable operating point. To reach this state required a substantial, dedicated effort from the US magnet team. The third attachment shows the strip-chart of the cool down; you can see where the upstream and downstream solenoids reach their operating temperature.

The cool-down of the focus-coil continues to go to plan. The magnet is now at around 80K and is expected to reach liquid-helium temperature at the end of this week.

Over the past few weeks substantial progress has been made towards completing the electrical and control systems for the magnets. This work culminated at the weekend with the start of quench-detector-system tests. The hard work and close coordination of Pierrick, the US magnet team and the DL electrical and controls teams have made this possible.

Commissioning and operations in ISIS Cycle 2015/01b

The electrical and controls experts are now preparing the systems for the start of magnet training. The goal is that magnet training will start in the week of the 20Jul15. In the run up to training, many systems need to be checked out and a formal “sign-off” procedure is being developed.

Tomorrow, TOF2, KL and EMR will be moved into their final Step IV positions. This will necessitate a further round of TOF calibrations. Steve Boyd, in consultation with the Analysis group and others, has developed a run plan. Data taking must be scheduled around magnet commissioning, so, items on the run plan will be executed in priority order as time permits.

It is important for the development of the experiment that the training of the MICE magnets is expedited. Therefore, in the coming days, preparations for magnet training, and then the magnet training itself, will have priority. Since we are setting up a new system and then training the magnets together for the first time, this period will necessarily be one in which plans are fluid and subject to change. I would therefore ask the shift crews to bear with, and support, Steve Boyd and Ryan Bayes (the MOM) as they respond to the evolving needs of the experiment.

This is an exciting time for MICE; are close to bringing Step IV into operation


Comments